OpenStudy (anonymous):

The probability that it will snow on sunday is 3/5. The probability that it will snow on both sunday and monday is 3/10. What is the probability that it will snow on monday, if it snowed on sunday.

OpenStudy (anonymous):

wouldn't the answer be 3/10? it seems it was given in the problem, unless i am reading it wrong >probability of snowing on both Sun and Mon is 3/10 >What is the probability of it snowing on Mon, it it snowed on Sunday? (thus, what is the probability of it snowing on both days?) it must be 3/10

OpenStudy (lgbasallote):

HINT!!! probability that it will snow on BOTH sunday AND monday is probability it will snow on sunday x probability it will snow on monday let \(P_s\) = probability it will snow on sunday \(P_m\) probability it will snow on monday \(P_{ms}\) probability it will snow on both dnays \[P_{ms} = P_s \times P_m\] \[\frac{P_{ms}}{P_s} = P_m\] \[\frac{\frac{3}{10}}{\frac{3}{5}} = P_m\] complex fractions... \[\frac{3}{10} \times \frac{5}{3} = P_m\]

OpenStudy (lgbasallote):

do you get it @Tony93 ?

OpenStudy (lgbasallote):

unless i understood it wrong o.O it's one way or the other lol

OpenStudy (lgbasallote):

question isnt really that clear

OpenStudy (anonymous):

yeah that is a confusing question; it's probably asking for the probability of snow falling on Monday, if it was snowing on Sunday, but then that means it was snowing on both days, which means the answer must be 3/10, as the problem states the probability of that happening is 3/10, so any other answer is illogical - it's like a loophole in the problem lol